Factory Overhead Cost Variances

Blumen Textiles Corporation began April with a budget for 37,000 hours of production in the Weaving Department. The department has a full capacity of 49,000 hours under normal business conditions. The budgeted overhead at the planned volumes at the beginning of April was as follows:

Variable overhead $99,900
Fixed overhead 68,600
Total $168,500

The actual factory overhead was $170,500 for April. The actual fixed factory overhead was as budgeted. During April, the Weaving Department had standard hours at actual production volume of 38,000 hours. Determine the variable factory overhead controllable variance and the fixed factory overhead volume variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Round your interim computations to the nearest cent, if required.

a. Variable factory overhead controllable variance: $  Favorable

b. Fixed factory overhead volume variance: $ Unfavorable
Answer #1

Solution a:

Standard variable overhead rate = $99,900 / 37000 = $2.70 per hour

Variable factory overhead controllable variance = Actual variable overhead - Standard variable overhead

= ($170,500 - $68,600) - (38000*$2.70)

= - $700 Favorable

Solution b:

Fixed overhead rate = $68,600 / 49000 = $1.40 per hour

Fixed overhead volume variance = Budgeted fixed overhead - Overhead applied = $68,600 - (38000*$1.40) = $15,400 Unfavorable